$SPAD/src/input gonshor.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool gonshor.output
)set message test on
)set message auto off
)clear all

———-

2 Some examples of algebras in genetics

Literature: [WB] A. Woerz-Busekros: Algebras in Genetics, LNB 36, Springer-Verlag, Berlin etc. 1980.

2.1 Commutative, non-associative algebras

A Gonshor genetic algebra ([WB], p. 41-42) of dimension 4:

The coefficient ring: — * —

--S 1 of 98
R := FRAC POLY INT
--R
--R
--R   (1)  Fraction(Polynomial(Integer))
--R                                                                 Type: Domain
--E 1

———-

The following multiplication constants may be chosen arbitrarily (notice that we write ckij for c_{(}i, j)^{k}):

— * —

--S 2 of 98
(c100, c101, _
c200, c201, c202, c211, _
c300, c301, c302, c303, c311, c312, c322) : R
--R

--R                                                                   Type: Void
--E 2
--S 3 of 98
c100 :=  1 ;     c101 := -1 ;
--R
--R
--R                                          Type: Fraction(Polynomial(Integer))
--E 3
--S 4 of 98
c200 :=  0 ;     c201 :=  1 ;     c202 := -1 ;
                 c211 :=  2 ;
--R
--R
--R                                          Type: Fraction(Polynomial(Integer))
--E 4
--S 5 of 98
c300 :=  1 ;     c301 :=  0 ;     c302 := -1 ;     c303 :=  1 ;
                 c311 :=  1 ;     c312 :=  0 ;
                                  c322 :=  2 ;
--R
--R
--R                                          Type: Fraction(Polynomial(Integer))
--E 5

———-

The matrices of the multiplication constants:

— * —

--S 6 of 98
gonshor : List SquareMatrix(4,R) :=
  [matrix [ [1, 0, 0, 0], [0, 0, 0, 0],_
            [0, 0, 0, 0], [0, 0, 0, 0] ],_
   matrix [ [c100, c101, 0, 0], [c101, 0, 0, 0],_
            [0, 0, 0, 0], [0, 0, 0, 0] ],_
   matrix [ [c200, c201, c202, 0], [c201, c211, 0, 0],_
            [c202, 0, 0, 0], [0, 0, 0, 0] ],_
   matrix [ [c300, c301, c302, c303], [c301, c311, c312, 0],_
            [c302, c312, c322, 0], [c303, 0, 0, 0] ] ] ;
--R
--R
--R                    Type: List(SquareMatrix(4,Fraction(Polynomial(Integer))))
--E 6
--S 7 of 98
basisSymbols : List Symbol := [subscript(e,[i]) for i in 0..3]
--R

--R
--R   (7)  [e ,e ,e ,e ]
--R          0  1  2  3
--R                                                           Type: List(Symbol)
--E 7
--S 8 of 98
GonshorGenetic := ALGSC(R, 4, basisSymbols, gonshor)
--R
--R
--R   (8)
--R  AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e
--R  1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--R                                                                 Type: Domain
--E 8
--S 9 of 98
commutative?()$GonshorGenetic
--R
--R   algebra is commutative
--R
--R   (9)  true
--R                                                                Type: Boolean
--E 9
--S 10 of 98
associative?()$GonshorGenetic
--R
--R   algebra is not associative
--R
--R   (10)  false
--R                                                                Type: Boolean
--E 10

———-

The canonical basis:

— * —

--S 11 of 98
e0 : GonshorGenetic := [1, 0, 0, 0] :: Vector R ;
--R
--R
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 11
--S 12 of 98
e1 : GonshorGenetic := [0, 1, 0, 0] :: Vector R ;
--R
--R

--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 12
--S 13 of 98
e2 : GonshorGenetic := [0, 0, 1, 0] :: Vector R ;
--R
--R
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 13
--S 14 of 98
e3 : GonshorGenetic := [0, 0, 0, 1] :: Vector R ;
--R
--R
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 14

———-

A generic element of the algebra:

— * —

--S 15 of 98
x  : GonshorGenetic := x0*e0 + x1*e1 + x2*e2 + x3*e3
--R
--R
--R   (15)  x3 e  + x2 e  + x1 e  + x0 e
--R             3       2       1       0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 15

———-

The matrix of the left multiplication with x :

— * —

--S 16 of 98
Lx := leftRegularRepresentation x
--R
--R
--R         +x0  - x1 + x0  - x2 + x1  x3 - x2 + x0+
--R         |                                      |
--R         |0     - x0     2x1 + x0        x1     |
--R   (16)  |                                      |
--R         |0       0        - x0       2x2 - x0  |
--R         |                                      |
--R         +0       0          0           x0     +
--R                                  Type: Matrix(Fraction(Polynomial(Integer)))
--E 16

———-

leftRegularRepresentationt 8 : GonshorGenetic -¿ R be the weight homomorphism defined by 8(e0) := 1 and 8(ei) := 0 for i = 1,2,3 . The coefficients of the characteristic polynomial of Lx depend only on 8(x) = x0 :

— * —

--S 17 of 98
p := characteristicPolynomial(Lx,Y)
--R
--R
--R           4     2  2    4
--R   (17)  x0  - 2Y x0  + Y
--R                                                    Type: Polynomial(Integer)
--E 17

———-

The left minimal polynomial of x divides Y * p(Y) :

— * —

--S 18 of 98
leftMinimalPolynomial x
--R
--R
--R          5      2 3     4
--R   (18)  ?  - 2x0 ?  + x0 ?
--R              Type: SparseUnivariatePolynomial(Fraction(Polynomial(Integer)))
--E 18
)clear prop A a b c r s
--S 19 of 98
A := GonshorGenetic
--R
--R
--R   (19)
--R  AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e
--R  1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--R                                                                 Type: Domain
--E 19
--S 20 of 98
a := x
--R
--R
--R   (20)  x3 e  + x2 e  + x1 e  + x0 e
--R             3       2       1       0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 20

--S 21 of 98
b := (1/4)*e1 + (1/5)*e2 + (3/20)*e3 + (2/5)*e0
--R
--R
--R          3      1      1      2
--R   (21)  -- e  + - e  + - e  + - e
--R         20  3   5  2   4  1   5  0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 21
--S 22 of 98
c := (1/3)*e1 + (1/7)*e2 + (8/21)*e3 + (1/7)*e0
--R
--R
--R          8      1      1      1
--R   (22)  -- e  + - e  + - e  + - e
--R         21  3   7  2   3  1   7  0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 22
--S 23 of 98
r  : R := r
--R
--R
--R   (23)  r
--R                                          Type: Fraction(Polynomial(Integer))
--E 23
--S 24 of 98
s  : R := s
--R
--R
--R   (24)  s
--R                                          Type: Fraction(Polynomial(Integer))
--E 24
--S 25 of 98
b*c
--R
--R
--R         2      1       47       2
--R   (25)  - e  + - e  - --- e  + -- e
--R         7  3   4  2   420  1   35  0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 25
--S 26 of 98
(b*c)*b
--R
--R

--R          893       277       4       4
--R   (26)  ---- e  - ---- e  + -- e  + --- e
--R         8400  3   1400  2   75  1   175  0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 26
--S 27 of 98
b*(c*b)
--R
--R
--R          893       277       4       4
--R   (27)  ---- e  - ---- e  + -- e  + --- e
--R         8400  3   1400  2   75  1   175  0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 27

———-

A: Algebra a,b,c : A r,s : R — * —

)clear prop AP
--S 28 of 98
AP := ALGPKG(R,A)
--R
--R
--R   (28)
--R  AlgebraPackage(Fraction(Polynomial(Integer)),AlgebraGivenByStructuralConstant
--R  s(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MA
--R  TRIX,MATRIX]))
--R                                                                 Type: Domain
--E 28
--S 29 of 98
r*a
--R
--R
--R   (29)  r x3 e  + r x2 e  + r x1 e  + r x0 e
--R               3         2         1         0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 29
--S 30 of 98
a*r
--R
--R
--R   (30)  r x3 e  + r x2 e  + r x1 e  + r x0 e
--R               3         2         1         0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])

--E 30
--S 31 of 98
a*b
--R
--R
--R         8x3 + 5x1 + 7x0      - 8x2 + 18x1 + x0      - 8x1 + 3x0      2x0
--R   (31)  --------------- e  + ----------------- e  + ----------- e  + --- e
--R                20        3           20         2        20      1    5   0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 31
--S 32 of 98
b*c
--R
--R
--R         2      1       47       2
--R   (32)  - e  + - e  - --- e  + -- e
--R         7  3   4  2   420  1   35  0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 32
--S 33 of 98
12 * c
--R
--R
--R         32      12            12
--R   (33)  -- e  + -- e  + 4e  + -- e
--R          7  3    7  2     1    7  0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 32
--S 34 of 98
(-3) * a
--R
--R
--R   (34)  - 3x3 e  - 3x2 e  - 3x1 e  - 3x0 e
--R                3        2        1        0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 34
--S 35 of 98
d  :=  a ** 12
--R
--R
--R   (35)
--R             11        10  2         9  2        10         11           8  4
--R         12x0  x3 + 4x0  x2  + (144x0 x1  + 144x0  x1 - 68x0  )x2 + 248x0 x1
--R       +
--R                9  3       10  2        11         12

--R         - 784x0 x1  - 86x0  x1  + 204x0  x1 - 24x0
--R    *
--R       e
--R        3
--R   +
--R         11         10  2       11            11       12        12
--R     (4x0  x2 - 92x0  x1  + 28x0  x1)e  + (4x0  x1 - x0  )e  + x0  e
--R                                      2                    1        0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 35
--S 36 of 98
-d
--R
--R
--R   (36)
--R               11        10  2           9  2        10         11
--R         - 12x0  x3 - 4x0  x2  + (- 144x0 x1  - 144x0  x1 + 68x0  )x2
--R       +
--R                8  4        9  3       10  2        11         12
--R         - 248x0 x1  + 784x0 x1  + 86x0  x1  - 204x0  x1 + 24x0
--R    *
--R       e
--R        3
--R   +
--R           11         10  2       11              11       12        12
--R     (- 4x0  x2 + 92x0  x1  - 28x0  x1)e  + (- 4x0  x1 + x0  )e  - x0  e
--R                                        2                      1        0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 36
--S 37 of 98
a + b
--R
--R
--R         20x3 + 3      5x2 + 1      4x1 + 1      5x0 + 2
--R   (37)  -------- e  + ------- e  + ------- e  + ------- e
--R            20     3      5     2      4     1      5     0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 37
--S 38 of 98
d-c
--R
--R
--R   (38)
--R                11         10  2          9  2         10           11
--R           252x0  x3 + 84x0  x2  + (3024x0 x1  + 3024x0  x1 - 1428x0  )x2
--R         +
--R                 8  4          9  3         10  2         11          12

--R           5208x0 x1  - 16464x0 x1  - 1806x0  x1  + 4284x0  x1 - 504x0   - 8
--R      /
--R         21
--R    *
--R       e
--R        3
--R   +
--R         11          10  2        11                11        12
--R     28x0  x2 - 644x0  x1  + 196x0  x1 - 1      12x0  x1 - 3x0   - 1
--R     ------------------------------------- e  + -------------------- e
--R                       7                    2             3           1
--R   +
--R        12
--R     7x0   - 1
--R     --------- e
--R         7      0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 38
--S 39 of 98
(a*(a*a) = leftPower(a,3)) :: Boolean
--R
--R
--R   (39)  true
--R                                                                Type: Boolean
--E 39
--S 40 of 98
(a ** 11 =  (a**8 * a**2) * a) :: Boolean
--R
--R
--R   (40)  true
--R                                                                Type: Boolean
--E 40
--S 41 of 98
(a ** 11 =  a**8 * (a**2 * a)) :: Boolean
--R
--R
--R   (41)  false
--R                                                                Type: Boolean
--E 41
--S 42 of 98
zero := 0$A
--R
--R
--R   (42)  0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 42

--S 43 of 98
zero : A := 0
--R
--R
--R   (43)  0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 43
--S 44 of 98
alternative?()$A
--R
--R   algebra is not left alternative
--R
--R   (44)  false
--R                                                                Type: Boolean
--E 44
--S 45 of 98
antiCommutative?()$A
--R
--R   algebra is not anti-commutative
--R
--R   (45)  false
--R                                                                Type: Boolean
--E 45
--S 46 of 98
associative?()$A
--R
--R   algebra is not associative
--R
--R   (46)  false
--R                                                                Type: Boolean
--E 46
--S 47 of 98
commutative?()$A
--R
--R   algebra is commutative
--R
--R   (47)  true
--R                                                                Type: Boolean
--E 47
--S 48 of 98
commutator(a,b)
--R
--R
--R   (48)  0

--S 49 of 98
antiCommutator(a,b)
--R
--R
--R         8x3 + 5x1 + 7x0      - 8x2 + 18x1 + x0      - 8x1 + 3x0      4x0
--R   (49)  --------------- e  + ----------------- e  + ----------- e  + --- e
--R                10        3           10         2        10      1    5   0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 49
--S 50 of 98
associator(a,b,c)
--R
--R
--R         - 21x2 + 6x1 + 7x0      12x2 - 30x1 + 58x0      12x1 - 28x0
--R   (50)  ------------------ e  + ------------------ e  + ----------- e
--R                 42          3           105         2       105      1
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 50
--S 51 of 98
basis()$A
--R
--R
--R   (51)  [e ,e ,e ,e ]
--R           0  1  2  3
--RType: Vector(AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX]))
--E 51
--S 52 of 98
n := rank()$A
--R
--R
--R   (52)  4
--R                                                        Type: PositiveInteger
--E 52
--S 53 of 98
v : Vector R := [i for i in 1..n]
--R
--R
--R   (53)  [1,2,3,4]
--R                                  Type: Vector(Fraction(Polynomial(Integer)))
--E 53
--S 54 of 98
g : A := represents  v

--R
--R
--R   (54)  4e  + 3e  + 2e  + e
--R           3     2     1    0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 54
--S 55 of 98
coordinates a
--R
--R
--R   (55)  [x0,x1,x2,x3]
--R                                  Type: Vector(Fraction(Polynomial(Integer)))
--E 55
--S 56 of 98
coordinates [a,b]
--R
--R
--R         +x0  x1  x2  x3+
--R         |              |
--R   (56)  |2   1   1    3|
--R         |-   -   -   --|
--R         +5   4   5   20+
--R                                  Type: Matrix(Fraction(Polynomial(Integer)))
--E 56
--S 57 of 98
a.3
--R
--R
--R   (57)  x2
--R                                          Type: Fraction(Polynomial(Integer))
--E 57
--S 58 of 98
flexible?()$A
--R
--R   algebra is flexible
--R
--R   (58)  true
--R                                                                Type: Boolean
--E 58
--S 59 of 98
leftAlternative?()$A
--R
--R   algebra is not left alternative
--R
--R   (59)  false

--R                                                                Type: Boolean
--E 59
--S 60 of 98
rightAlternative?()$A
--R
--R   algebra is not right alternative
--R
--R   (60)  false
--R                                                                Type: Boolean
--E 60
--S 61 of 98
sB := someBasis()$A
--R
--R
--R   (61)  [e ,e ,e ,e ]
--R           0  1  2  3
--RType: Vector(AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX]))
--E 61
--S 62 of 98
zero? a
--R
--R
--R   (62)  false
--R                                                                Type: Boolean
--E 62
--S 63 of 98
associatorDependence()$A
--R
--R
--R   (63)  [[1,1,1,0,0,0],[0,1,0,1,0,0],[1,0,0,0,1,0],[- 1,- 1,0,0,0,1]]
--R                            Type: List(Vector(Fraction(Polynomial(Integer))))
--E 63

———-

ConditionsForIdempotents()$A — * —

--S 64 of 98
jacobiIdentity?()$A
--R
--R   Jacobi identity does not hold
--R
--R   (64)  false
--R                                                                Type: Boolean
--E 64

--S 65 of 98
jordanAlgebra?()$A
--R
--R   algebra is commutative
--R   this is not a Jordan algebra
--R
--R   (65)  false
--R                                                                Type: Boolean
--E 65
--S 66 of 98
jordanAdmissible?()$A
--R
--R   algebra is not Jordan admissible
--R
--R   (66)  false
--R                                                                Type: Boolean
--E 66
--S 67 of 98
lieAdmissible?()$A
--R
--R   algebra is Lie admissible
--R
--R   (67)  true
--R                                                                Type: Boolean
--E 67

———-

ConditionsForIdempotents — * —

--S 68 of 98
b2 := [reduce(+,[sB.i for i in 1..k]) for k in 1..n]
--R
--R
--R   (68)  [e ,e  + e ,e  + e  + e ,e  + e  + e  + e ]
--R           0  1    0  2    1    0  3    2    1    0
--RType: List(AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX]))
--E 68
--S 69 of 98
coordinates  (a ,b2 :: Vector A)
--R
--R
--R   (69)  [- x1 + x0,- x2 + x1,- x3 + x2,x3]
--R                                  Type: Vector(Fraction(Polynomial(Integer)))
--E 69

--S 70 of 98
coordinates  ([a,b] ,bb := (b2 :: Vector A))
--R
--R
--R         +- x1 + x0  - x2 + x1  - x3 + x2  x3+
--R         |                                   |
--R   (70)  |    3          1          1       3|
--R         |   --         --         --      --|
--R         +   20         20         20      20+
--R                                  Type: Matrix(Fraction(Polynomial(Integer)))
--E 70
--S 71 of 98
leftMinimalPolynomial a
--R
--R
--R          5      2 3     4
--R   (71)  ?  - 2x0 ?  + x0 ?
--R              Type: SparseUnivariatePolynomial(Fraction(Polynomial(Integer)))
--E 71
--S 72 of 98
leftPower (a,10)
--R
--R
--R   (72)
--R          9        8  2          7  2      8        9          8  2      10
--R     (10x0 x3 - 6x0 x2  + (- 32x0 x1  + 8x0 x1 + 2x0 )x2 + 13x0 x1  + 5x0  )e
--R                                                                             3
--R   +
--R           9         8  2      9        10            9       10        10
--R     (- 2x0 x2 + 26x0 x1  + 6x0 x1 - 4x0  )e  + (- 2x0 x1 + x0  )e  + x0  e
--R                                            2                     1        0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 72
--S 73 of 98
rightPower(a,10)
--R
--R
--R   (73)
--R          9        8  2          7  2      8        9          8  2      10
--R     (10x0 x3 - 6x0 x2  + (- 32x0 x1  + 8x0 x1 + 2x0 )x2 + 13x0 x1  + 5x0  )e
--R                                                                             3
--R   +
--R           9         8  2      9        10            9       10        10
--R     (- 2x0 x2 + 26x0 x1  + 6x0 x1 - 4x0  )e  + (- 2x0 x1 + x0  )e  + x0  e
--R                                            2                     1        0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])

--E 73
--S 74 of 98
leftRegularRepresentation a
--R
--R
--R         +x0  - x1 + x0  - x2 + x1  x3 - x2 + x0+
--R         |                                      |
--R         |0     - x0     2x1 + x0        x1     |
--R   (74)  |                                      |
--R         |0       0        - x0       2x2 - x0  |
--R         |                                      |
--R         +0       0          0           x0     +
--R                                  Type: Matrix(Fraction(Polynomial(Integer)))
--E 74
--S 75 of 98
leftRegularRepresentation (a,bb)
--R
--R
--R         +  x1     x2 - 2x1 + x0     - x3 + x1 - x0        x3 - x2 + x0   +
--R         |                                                                |
--R         |x1 + x0  x2 - 4x1 - x0       - x3 + 2x1        x3 - x2 + x1 + x0|
--R   (75)  |                                                                |
--R         |x1 + x0    x2 - 4x1       - x3 - 2x2 + 2x1       x3 + x2 + x1   |
--R         |                                                                |
--R         +x1 + x0    x2 - 4x1     - x3 - 2x2 + 2x1 - x0  x3 + x2 + x1 + x0+
--R                                  Type: Matrix(Fraction(Polynomial(Integer)))
--E 75
--S 76 of 98
leftUnit()$A
--R
--R   this algebra has no left unit
--R
--R   (76)  "failed"
--R                                                    Type: Union("failed",...)
--E 76
--S 77 of 98
represents (v,bb)
--R
--R
--R   (77)  4e  + 7e  + 9e  + 10e
--R           3     2     1      0
--RType: AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX])
--E 77
--S 78 of 98
rightMinimalPolynomial a

--R
--R
--R          5      2 3     4
--R   (78)  ?  - 2x0 ?  + x0 ?
--R              Type: SparseUnivariatePolynomial(Fraction(Polynomial(Integer)))
--E 78
--S 79 of 98
rightRegularRepresentation a
--R
--R
--R         +x0  - x1 + x0  - x2 + x1  x3 - x2 + x0+
--R         |                                      |
--R         |0     - x0     2x1 + x0        x1     |
--R   (79)  |                                      |
--R         |0       0        - x0       2x2 - x0  |
--R         |                                      |
--R         +0       0          0           x0     +
--R                                  Type: Matrix(Fraction(Polynomial(Integer)))
--E 79
--S 80 of 98
rightRegularRepresentation (a,bb)
--R
--R
--R         +  x1     x2 - 2x1 + x0     - x3 + x1 - x0        x3 - x2 + x0   +
--R         |                                                                |
--R         |x1 + x0  x2 - 4x1 - x0       - x3 + 2x1        x3 - x2 + x1 + x0|
--R   (80)  |                                                                |
--R         |x1 + x0    x2 - 4x1       - x3 - 2x2 + 2x1       x3 + x2 + x1   |
--R         |                                                                |
--R         +x1 + x0    x2 - 4x1     - x3 - 2x2 + 2x1 - x0  x3 + x2 + x1 + x0+
--R                                  Type: Matrix(Fraction(Polynomial(Integer)))
--E 80
--S 81 of 98
rightUnit()$A
--R
--R   this algebra has no right unit
--R
--R   (81)  "failed"
--R                                                    Type: Union("failed",...)
--E 81
--S 82 of 98
structuralConstants()$A
--R
--R
--R          +1  0  0  0+ + 1   - 1  0  0+ + 0   1  - 1  0+ + 1   0  - 1  1+
--R          |          | |              | |              | |              |

--R          |0  0  0  0| |- 1   0   0  0| | 1   2   0   0| | 0   1   0   0|
--R   (82)  [|          |,|              |,|              |,|              |]
--R          |0  0  0  0| | 0    0   0  0| |- 1  0   0   0| |- 1  0   2   0|
--R          |          | |              | |              | |              |
--R          +0  0  0  0+ + 0    0   0  0+ + 0   0   0   0+ + 1   0   0   0+
--R                          Type: Vector(Matrix(Fraction(Polynomial(Integer))))
--E 82
--S 83 of 98
structuralConstants(bb)
--R
--R
--R          +0  1  1  1+ + 1   - 1   0    0 + +- 1  0   0   - 1+ +1  1  0  1+
--R          |          | |                  | |                | |          |
--R          |1  2  2  2| |- 1  - 5  - 4  - 4| | 0   2   2    1 | |1  2  1  2|
--R   (83)  [|          |,|                  |,|                |,|          |]
--R          |1  2  2  2| | 0   - 4  - 3  - 3| | 0   2   0   - 1| |0  1  2  3|
--R          |          | |                  | |                | |          |
--R          +1  2  2  2+ + 0   - 4  - 3  - 3+ +- 1  1  - 1  - 2+ +1  2  3  4+
--R                          Type: Vector(Matrix(Fraction(Polynomial(Integer))))
--E 83
--S 84 of 98
unit()$A
--R
--R   this algebra has no unit
--R
--R   (84)  "failed"
--R                                                    Type: Union("failed",...)
--E 84
--S 85 of 98
biRank  a
--R
--R
--R   (85)  4
--R                                                        Type: PositiveInteger
--E 85
--S 86 of 98
leftRank a
--R
--R
--R   (86)  4
--R                                                        Type: PositiveInteger
--E 86
--S 87 of 98
doubleRank a
--R

--R
--R   (87)  4
--R                                                        Type: PositiveInteger
--E 87
--S 88 of 98
rightRank a
--R
--R
--R   (88)  4
--R                                                        Type: PositiveInteger
--E 88
--S 89 of 98
weakBiRank a
--R
--R
--R   (89)  4
--R                                                        Type: PositiveInteger
--E 89
--S 90 of 98
basisOfCenter()$AP
--R
--R
--R   (90)  [e ]
--R           3
--RType: List(AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX]))
--E 90
--S 91 of 98
basisOfLeftNucleus()$AP
--R
--R
--R   (91)  [e ]
--R           3
--RType: List(AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX]))
--E 91
--S 92 of 98
basisOfNucleus()$AP
--R
--R
--R   (92)  [e ]
--R           3
--RType: List(AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX]))
--E 92
--S 93 of 98
basisOfRightNucleus()$AP

--R
--R
--R   (93)  [e ]
--R           3
--RType: List(AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX]))
--E 93
--S 94 of 98
basisOfCentroid()$AP
--R
--R
--R          +0  0  0  0+ +1  0  0  0+
--R          |          | |          |
--R          |0  0  0  0| |0  1  0  0|
--R   (94)  [|          |,|          |]
--R          |0  0  0  0| |0  0  1  0|
--R          |          | |          |
--R          +1  0  0  0+ +0  0  0  1+
--R                            Type: List(Matrix(Fraction(Polynomial(Integer))))
--E 94
--S 95 of 98
basisOfCommutingElements()$AP
--R
--R
--R   (95)  [e ,e ,e ,e ]
--R           3  2  1  0
--RType: List(AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX]))
--E 95
--S 96 of 98
basisOfLeftNucloid()$AP
--R
--R
--R          +0  0  0  0+ +1  0  0  0+
--R          |          | |          |
--R          |0  0  0  0| |0  1  0  0|
--R   (96)  [|          |,|          |]
--R          |0  0  0  0| |0  0  1  0|
--R          |          | |          |
--R          +1  0  0  0+ +0  0  0  1+
--R                            Type: List(Matrix(Fraction(Polynomial(Integer))))
--E 96
--S 97 of 98
basisOfMiddleNucleus()$AP
--R
--R
--R   (97)  [e ]
--R           3

--RType: List(AlgebraGivenByStructuralConstants(Fraction(Polynomial(Integer)),4,[*01e0,*01e1,*01e2,*01e3],[MATRIX,MATRIX,MATRIX,MATRIX]))
--E 97
--S 98 of 98
basisOfRightNucloid()$AP
--R
--R
--R          +0  0  0  0+ +1  0  0  0+
--R          |          | |          |
--R          |0  0  0  0| |0  1  0  0|
--R   (98)  [|          |,|          |]
--R          |0  0  0  0| |0  0  1  0|
--R          |          | |          |
--R          +1  0  0  0+ +0  0  0  1+
--R                            Type: List(Matrix(Fraction(Polynomial(Integer))))
--E 98
)spool
)lisp (bye)

———-

References^{[1] nothing}

25